A firm can sell 10 gizmos when the market price is £40 and can sell 60 gizmos when the market price is £15. The firm faces constant marginal cost of £15, with no other costs of production a) Assuming demand is linear, find the market demand function for gizmos and the associated inverse demand function. b) Determine the price elasticity of demand when the market price is £40. c) Suppose the firm is a monopolist in this market. i.) Explain the connection between marginal revenue and price elasticity of L) Explain the connection between price elasticity of demand and the firms monopoly power Calculate the firms profit maximising output. Quantify the firms monopoly power. iii.) d) Now suppose there are two identical firms competing in this market, characterised by the above demand and costs. Solve for the Cournot equilibrium outputs, market price and profit levels.
Answer #1

1A.b. p1 = 40

q1 = 10

p2 = 15

q2 = 60

MC = 15

price elasticity = 60-10 X 40

   15-40    10

   = (50/35) X (40/10)

   = 5.71

c.1 MR = P (1- !/e)

2. Monopoly power increases with lower elastic demand and decrease with higher elastic demand